OpenStudy (anonymous):

What type is the curve and sketch the graph. Do not plot points. 4x^2 + y^2 = 1 I know it's an ellipse, but how do I tell how tall and how wide it should be??

OpenStudy (anonymous):

Compare to the general formula for an ellipse: \[\left(\frac{x}{a}\right)^2+\left(\frac{y}{b}\right)^2=1\] \(a\) is the length of the horizontal "radius" and \(b\) is the length of the vertical "radius".

OpenStudy (anonymous):

Should I do anything with the 4?

OpenStudy (anonymous):

so it would be x= 1/4 and -1/4 and y = 1, -1. for the edges of the ellipse?

OpenStudy (anonymous):

My book is saying that it should be 1/2 instead of 1/4th?

OpenStudy (anonymous):

Sorry, I mean \(a^2=\dfrac{1}{4}\), which means \(a=\dfrac{1}{2}\).

OpenStudy (anonymous):

OH, perfect sense now

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\\ -------------\\ 4x^2 + y^2 = 1 \implies \cfrac{(x-0)^2}{\frac{1}{4}}+\cfrac{(y-0)^2}{1} = 1\)